# Conceptualization, Data Curation, Software, Visualization, Writing - original draft

### 01 Mar 2026 v5

@@ 4.1.3 Scenario Generation for CHA-UC
Figure 6: (Case Study I) Annual scatter plot of temperatures of line 192 (Line with biggest temperature difference) obtained through post-hoc evaluation. 20-day moving average (MA) lines are drawn above. Meteorogical seasons are colored in the background. (a) The temperatures of DLR, SLR and CHA-UC are compared. A zoom-in of thermal overload temperatures during periods when the MA of SLR is higher than CHA-UC, is provided. (b) The temperatures of QRF and CHA-UC are compared.
4.1.4 Benchmark Model
A quantile regression forest (QRF) model was used to benchmark the performance of the CHA-UC model. QRF is an ensemble method that generates multiple decision trees from the training data and produces quantile forecasts from the forest output. The model setting was adopted from Dupin et al. Dupin2019. The forest was created using 1-year data from 2022. The confidence level was set to 90%, and the overestimation frequency was 7.4%. The QRF forecasts are conservative and capture the overall temporal pattern of the actual DLR. The quantile trajectories move with the observations, although sharp peaks are smoothed as a result of the ensemble-based quantile estimation. For the QRF-UC simulation, the DLR line constraints (11h) were replaced with QRF forecasts.
4.2 Case study I - Value of Conductor Health Awareness
In case study I, wind forecast errors are neglected ($\hat{P}^{\text{wind}}_{w,t}=P^{\text{wind}}_{w,t}$) and only DLR forecast errors are considered. This simplifies the risk structure and reduces the disparity between the day-ahead operation and the real-time operation, allowing for a clearer interpretation of the results. Also, the minimum online up-reserve requirement is set to zero, as this case study does not account for WP forecast errors.
4.2.1 Overall Performance Comparison
Table 1 presents a breakdown of the total cost of four models: baseline SLR, baseline DLR, QRF and CHA-UC. The units are normalized to one day. The day-ahead cost corresponds to the objective value of (9), which includes UC costs, fuel costs and the day-ahead load shedding costs. The total cost is the sum of day-ahead costs, real-time costs and conductor depreciation costs. The benchmark risk-averse measure, QRF-UC does not show advantage over baseline SLR, since its lower DA costs (21.29 vs. 21.44) are offset by the higher depreciation cost (0.31 vs. 0.16). Simulations with higher confidence level (95%) shows even higher total costs. CHA-UC achieves the lowest annual total cost of 21.44 M\$/day, reducing costs by 2.05%, 0.74% and 0.74% compared to DLR, SLR and QRF, respectively. This improvement reflects its ability to balance the risk of ETO against reduced operation costs. Although CHA-UC adopts DLR, it applies a conservative line flow than the baseline by accounting for expected conductor depreciation cost. As a result, the DLR baseline yields the lowest day-ahead cost, followed by CHA-UC and then SLR. WP curtailments are reduced by 89.1% (9.1 v.s. 83.4) when DLR is used instead of SLR, as it relieves grid congestion. CHA-UC retains most of this advantage, though its robustness discards some of it, leading to a 85.0% reduction (12.5 v.s. 83.4). Using QRF retains the least of this advantage, leading to a 66.1% reduction (28.2 vs. 83.4). On the other hand, depreciation costs were lowest for CHA-UC followed by SLR and DLR indicating that the CHA-UC preserves conductor lifespan most effectively. Finally, reserves are only activated in CHA-UC, as the SLR and DLR baselines involve no uncertainty in this case study.
@@ 4.2.1 Overall Performance Comparison
4.2.2 Risk of ETO
@@
Figure 6(a) visualizes the annual line temperature from the three models. In spring, fall and winter, the line temperatures were high in order of DLR baseline, CHA-UC and SLR baseline. This order originates from DLR not considering stochasticity with high ratings, CHA-UC considering stochasticity with high ratings, and SLR using a uniform low rating. However, during summer, the overall line temperatures increase and the temperature of SLR exceeds CHA-UC. This is because the hot operating conditions of summer increases the risk of SLR. Since summer weathers are close to the conservative conditions for SLR calculation, the “conservative” SLR becomes risky as well. Thus, CHA-UC chooses a more conservative loading, shown in the day-ahead cost of 29.99 M\$/day, which is higher than 29.96 M\$/day of SLR baseline. A zoom-in of the scatter plot exhibits the results during this period. The temperatures from SLR was stamped more commonly over the thermal limit compared to CHA-UC, although CHA-UC has some outliers. This resulted in a CHA-UC depreciation cost of 0.13 M\$/day which was 3.5 times and 7.3 times less than SLR and DLR baseline, respectively.
@@
On the other hand, QRF chooses an even more conservative operation in summer, costing 30.35 M\$/day, reflected in Fig. 6(b). However, the cost measures were relatively ineffective, leading to a higher depreciation cost compared to CHA-UC (0.19 vs. 0.13). Despite the moving average temperature of QRF being similar with CHA-UC, the QRF outliers tend to show more common high temperature ETOs. The distribution in the zoom-in area shows that CHA-UC exploits the knowledge on the state of the conductor, balancing low-temperature ETOs with reduced operation costs, while QRF-UC frequently operates in higher temperature regions. This results in higher conductor depreciation costs of QRF albeit the higher expenditure for conservativeness.
4.2.3 Effect of CHA-UC to Commitment Decisions
@@
The annual average of commitment decisions made by DLR baseline is illustrated on a Texan map in Fig. 7(a). The cheapest units which comply with UC and DC PF constraints are committed, leading to aggressive line loading. The employment of CHA-UC leads to committing additional costly units to hedge the risks of ETO, leading to an overall relief of line loading in DLR lines, as presented in Fig. 7(b). For example, line 74, a key line connecting the south-west region and Dallas, is highly loaded in a direction to Dallas (Bus 22 to Bus 94) with an annual average of 84.5% (normalized by line capacity) when operated by DLR baseline. The transition to CHA-UC led to an 8.7%p loading relief. This contributes to an annual conductor health relief of 14.0%, which translates to a 112.4 M\$ reduction in depreciation costs. Here, the commitment decisions took a major role; the annual average commitment on bus 94 increased by 39%, and the commitment on bus 30, adjacent to the east of bus 94, increased by 34%, leading to the reduction of power flow to Dallas and the stress on line 74.
4.3 Case study II - Impact of Correlated Forecast Errors
@@
Figure 8: (Case Study II) Lineflow, temperature and depreciation cost of line 74 at day 107 from baseline SLR, DLR and CHA-UC. (a) the day-ahead lineflow. Day-ahead DLR values and the security margin are colored orange. (b) real-time lineflow, obtained as a result of the operation stage. (c) The resultant conductor line temperature in lines and depreciation costs in bars, obtained through post-hoc analysis.
4.3.1 Overall Performance Comparison
The costs of the annual simulation are loaded on the right-hand side of Table 1. Overall, the trend of the three models from case study I was preserved. CHA-UC recorded the lowest annual total cost, 25.3 M\$, contributing to a 34.2%, 25.5% and 26.1% reduction compared to DLR, SLR and QRF, respectively. However, unlike case study I, CHA-UC always incurred the highest day-ahead cost and it recorded the least reserve activation costs. This occurs because only CHA-UC accounts the stochasticity in WP forecasts in the day-ahead decision making.
4.3.2 Impact of WP forecast error
Figure 8 depicts how WP forecast errors affect line flow and conductor depreciation. In the day-ahead UC problem, the line flow is limited to stay under the security margin $K_{l,t}$ of DLR forecasts. CHA-UC keeps a conservative DA line flow compared to DLR baseline by considering the stochasticity of DLR and WP forecasts, presented in Fig. 8(a). When the actual values of WP and DLR take effect, the real-time operation model deploys reserve to keep the power balance, creating fluctuations in the line flow shown in Fig. 8(b).
In hour 13 and 19, the RT line flow of DLR baseline surpassed the security margin of the DA DLR forecast, due to generation adjustments to balance WP forecast errors. On the other hand, in hour 21, the actual DLR values were lower than the forecast margin, causing ETO for DLR baseline. In all ETO cases above, WP and DLR forecast errors combined unfavorably to amplify the depreciation. The line temperature and depreciation cost due to RT line flow in Fig. 8(c) visualize the impact of DLR and WP forecast errors. The depreciation cost incurred from hours 13 and 19 were 0.23 M\$ and 0.12 M\$ caused by ETOs of 98.1°C and 97.6°C, respectively. The relatively large difference in depreciation cost from a modest temperature difference comes from two ways: 1) the depreciation cost increases exponentially by temperature and 2) the diminishing depreciation cost shown in Fig. 2 reduces the sensitivity to ETO for hour 19 after the ETO in hour 13.
4.3.3 Impact of Correlated WP and DLR Forecast Error
Two pairs of adjacent wind farms and lines were selected for an analysis on inter-asset correlation. Table 2 lists the pairs; the two pairs have same signs of 1) correlation coefficients and 2) generation shift factor (GSF) of the WP bus to the line, but have different signs of average lineflow.
Table 2: (Case Study II) Pairs of correlated wind generators and lines studied. The correlation was captured using Spearman’s correlation coefficients and the average line flow using DLR baseline is given. GSF and line flow were obtained by setting the paired bus as the from-bus.
@@ 4.3.3 Impact of Correlated WP and DLR Forecast Error
Figure 9 illustrates the heatmap of line temperature under baseline DLR. Forecast errors are calculated by subtracting actual values from the forecast value, meaning that a larger forecast error represents a larger overestimation of the value. Since the GSF is positive in pair 1, the shortage of WP leads to a reduction of line flow. The negative line flow becomes smaller, leading positive WP forecast error (shortage) to stress the conductor thermal limit.
WP forecast errors are often hedged by the DLR forecast errors. The underestimation of DLR leads to an offset, where the WP shortage in the fourth quadrant is covered by the abundant line capacity. The first quadrant becomes risk-intensive due to the aggravation of DLR overestimation. Conversely, for pair 2, negative WP forecast errors (abundancy) stress the conductor thermal limit, leading the second quadrant to be risk-intensive.
Table 3: (Case Study II) Annual average CHA-UC day-ahead lineflows before and after removing correlation from the scenario set
@@
1 -5.09 -5.38 Decreased
2 8.06 6.76 Increased
The positive correlation of the pairs causes the forecast error to occur more often in the first and third quadrants. Therefore, pair 1 becomes risk-prone, as the risk is dense in the first quadrant, and pair 2 becomes risk-averse, as the risk is dense in the second quadrant. This leads to a critical conclusion: the risk of positive forecast error correlation between WP and DLR is amplified when the GSF and average lineflow have a different sign, and the risk is hedged when they have a same sign. The emergent property due to correlation should be considered for DLR operation.
An experiment regarding correlation proves that the CHA-UC is successfully reflecting the emergent effect of correlation. The scenario indices of the pairs were scrambled, changing the marginal minimally, so that the correlation of the scenarios becomes close to 0. This will lead to the inconsideration of correlation. Separate experiments done with two pairs, revealed in Table 3, show that the original (correlation-considering) CHA-UC has kept a less conservative line flow for the risk-prone pair 1 and a more conservative line flow for the risk-hedging pair 2. This asserts that the CHA-UC successfully handles the effect of correlation.
5 Conclusion
This paper presents CHA-UC, a stochastic unit commitment problem that considers line depreciation from ETO, effective for DLR operation. The CHA-UC optimization contains a conductor temperature estimation module and a risk-base conductor depreciation cost function to co-optimize the operation cost and depreciation cost. The cost of conductor depreciation is quantified through risk-base valuation. In result, case studies showcase a 0.74% and 85% reduction of total costs and WP curtailments compared to SLR operation, respectively. Moreover, it outperformed the baseline risk-averse measure, QRF-UC with a 0.74% total cost difference, while DLR operation without risk consideration showed highest costs due to the large conductor depreciation costs. The soundness of CHA-UC was proven by comparing annual line temperatures with the spatial commitment differences. A correlation analysis of DLR and WP forecast errors reveals the emergent properties from interactions of correlation, GSF and line flow. CHA-UC handles this properly. The statistics and analysis provided in the case studies will be important resources for DLR integration. The computational simplicity of the conductor temperature linear approximation will pay off. Despite some assumptions, our model proves to be most effective in the day-ahead operation with DLR, and provides meaningful materials for future grid research.
[…]
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request.

### 14 Apr 2026 v6

@@ IV Case Study
Figure 4: (Case Study I) Annual scatter plot of temperatures of line 192 (Line with biggest temperature difference) obtained through post-hoc evaluation. 20-day moving average (MA) lines are drawn above. (a) The temperatures of DLR, SLR and CHA-UC are compared. A zoom-in of thermal overload temperatures is provided. (b) The temperatures of QRF and CHA-UC are compared.
IV-A Preliminaries
IV-A1 Simulation Settings
In all case studies, a set of 21 lines was selected for DLR application [^2] Lines were screened by solving an annual DCOPF with SLR and selecting those with the highest dual values of the line-flow constraint.. Each 345 kV line was modeled as ACSR Finch, and ampacity was adjusted to fit the TX-123BT system through a line multiplier. All optimization models were implemented in Julia/JuMP [Lubin2023] and solved with Gurobi. The LoTS limit was set to 10% (Section II-C). Corrosivity factor $\gamma$ was obtained from mapping steel corrosivity zone C2 and under to 1, C3 and above to 1.17, obtained from [IEC2015, EngineeringDirector]. The line cost factor $B_{l}=$\$936/MVA-km was obtained by excluding tower and loss components from a 400 kV, 3190 MVA case study [Parsons2012]. The PDF modeling line failure $f(\cdot)$ was modeled as a normal distribution following $\mu=0.1,\sigma=0.05$ [Teh2019]. The conductor thermal limit was set to 95°C. The loading-temperature linear fit was sampled from 95 °C and 150 °C, and the piecewise depreciation cost function was sampled every 10 °C from 95 °C. Conductor health was initialized to fit a mean LoTS rate of 1% (a sensible rate from the results of [Adomah2000]) by creating synthetic historical ETO exposures from a normal distribution. The historical ETO exposure temperatures are passed on between different horizons to update the conductor state properly. For DLR security margin, line limits were constrained as $K_{l,t}\hat{\overline{F}}^{\mathrm{DLR}}_{l,t}=\max(0.8\hat{\overline{F}}^{\mathrm{DLR}}_{l,t},\,\underline{F}^{\mathrm{SLR}}_{l})$. The load profile of 2021 given in the TX-123BT test system was used. The value of lost load, $C^{\text{VoLL}}$, is assumed to be \$3500/MWh following MISO before 2025/05 [Kirschen2018]. Reserve activation costs followed [Dvorkin2025]: $C_{i}^{\mathrm{r}+}$ equals three times, and $C_{i}^{\mathrm{r}-}$ equals 0.5 times, the linear cost coefficient of generator $i$.
IV-A2 Data Collection
Weather data for 2021–2022 were obtained from NOAA’s HRRR model [Benjamin2016], with 3 km spatial and 1 h temporal resolution. Forecasts issued at 18:00 on the previous day were used for day-ahead inputs, and analysis data were treated as actual conditions. The 2022 data were used to generate stochastic scenarios, while the 2021 data were used for annual simulations. WP profiles were derived from wind speeds using a standard power curve [Giorsetto1983] with hub-height correction, assuming uniform turbine specifications. For each line, the hourly minimum DLR among intersecting $\times3$ km grid cells was selected. HBE parameters followed IEEE [IEEE738], with solar heat gain fixed at a worst-case Texas condition.
IV-A3 Scenario Generation for CHA-UC
To generate representative WP and DLR forecast-error scenarios while preserving their correlation, a daily trajectory dataset of 24-hour forecast errors is built. Historically extreme trajectories are first seeded, then a greedy max–min selection adds diverse trajectories. Scenario probabilities are then assigned using nearest-neighbor weighting with penalties on large deviations, yielding a scenario set that reflects the empirical error distribution. A total of 20 scenarios were generated. The statistic and correlation structure were well preserved: the mean, standard deviation and correlation varied by 0.7%, 8% and 28% (Empirical 0.037 vs. Scenario 0.052), respectively.
((a))
((b))
@@ IV-A3 Scenario Generation for CHA-UC
IV-A4 Benchmark Model
A quantile regression forest (QRF) model was employed to benchmark the performance of the CHA-UC model. QRF is an ensemble method that generates multiple decision trees from the training data and produces quantile forecasts from the forest output. The model setting was adopted from Dupin et al. [Dupin2019]. The forest was created using 1-year data from 2022. The confidence level was set to 90%, and the overestimation frequency was 7.4%. Overall, the quantile trajectories move with the observations, although sharp peaks are smoothed as a result of the ensemble-based quantile estimation. For the QRF-UC simulation, the DLR line constraints (14h) were replaced with QRF forecasts.
IV-B Case study I - Value of Conductor Health Awareness
In case study I, wind forecast errors are neglected ($\hat{P}^{\text{wind}}_{w,t}=P^{\text{wind}}_{w,t}$) and only DLR forecast errors are considered. This simplifies the risk structure and reduces the disparity between the day-ahead operation and the real-time operation, allowing a clearer interpretation of the results. Also, the minimum online up-reserve requirement is set to zero, as this case study does not account for WP forecast errors.
IV-B1 Overall Performance Comparison
Table I presents a breakdown of the normalized total cost of four models: baseline SLR, baseline DLR, QRF and CHA-UC. Day-ahead cost is the objective value of (12), and total cost includes day-ahead, real-time, and conductor depreciation costs. QRF-UC does not outperform baseline SLR: its lower day-ahead cost (21.29 vs. 21.44) is offset by higher depreciation cost (0.32 vs. 0.16); using a 95% confidence level further increases total cost. Although the DLR baseline yields the lowest day-ahead cost, followed by CHA-UC, QRF and SLR, CHA-UC achieves the lowest annual total cost of 21.44 M\$/day, reducing costs by 2.01%, 0.75% and 0.79% compared to DLR, SLR and QRF, respectively. This improvement reflects its ability to balance the risk of ETO against reduced operation costs under DLR. Relative to SLR, WP curtailment is reduced by 85.8% with DLR (11.7 vs. 82.6), 82.1% with CHA-UC (14.8 vs. 82.6), and 67.9% with QRF (26.5 vs. 82.6), showing that CHA-UC retains most of the congestion-relief of DLR. On the other hand, depreciation costs were lowest for CHA-UC followed by SLR and DLR indicating that the CHA-UC preserves conductor lifespan most effectively. Finally, reserves are only activated in CHA-UC, as other benchmark baselines does not involve uncertainty in Case Study I.
IV-B2 Risk of ETO
@@
Figure 4(a) visualizes the annual line temperature under the three models. In spring, fall and winter, the line temperatures were high in order of DLR baseline, CHA-UC and SLR baseline. This reflects the use of high ratings without stochastic concerns in DLR, high ratings with stochastic concerns in CHA-UC, and a uniformly low rating in SLR. However, during summer, the overall line temperatures increase and the temperature of SLR exceeds CHA-UC. Since summer weathers are close to the conservative conditions for SLR calculation, the “conservative” SLR becomes risky as well. Thus, CHA-UC chooses a similarly conservative loading, shown in the day-ahead cost of 29.96 M\$/day for both CHA-UC and SLR Baseline. A zoomed-in scatter plot confirms that SLR and DLR temperatures more frequently exceed the thermal limit than those of CHA-UC. This resulted in a CHA-UC depreciation cost of 0.12 M\$/day which was 3.8 times and 8.1 times less than SLR and DLR baseline, respectively.
@@
On the other hand, QRF chooses an even more conservative operation in summer, costing 30.35 M\$/day, reflected in Fig. 4(b). Yet this conservativeness was relatively ineffective, leading to a higher depreciation cost compared to CHA-UC (0.19 vs. 0.12). Despite the moving average of QRF and CHA-UC are similar, the QRF outliers show more frequent high temperature ETOs. The distribution in the zoom-in area shows that CHA-UC exploits the knowledge on the state of the conductor, balancing low-temperature ETOs with reduced operation costs, while QRF-UC frequently operates in higher temperature regions without conductor health-awareness. This is also reflected in the annual ETO statistics of all DLR-applied lines: the ETO frequency of QRF-UC is 1.2 times higher (2397 vs. 1934 hours) but the total depreciation cost is 2.6 times greater (0.32 vs. 0.12 M\$/day) due to the difference in average ETO temperatures (102.5 °C vs. 100.7 °C). This result highlights the value of conductor-awareness of CHA-UC; conductor depreciation costs are minute albeit the smaller expenditure for conservativeness compared to the QRF benchmark model.
IV-B3 Effect of CHA-UC to Commitment Decisions
@@
The annual average of commitment decisions made by DLR baseline is illustrated on a Texan map in Fig. 5(a). The cheapest units that comply with the UC and PF constraints are committed, leading to aggressive line loading. The employment of CHA-UC leads to committing additional costly units to hedge the risks of ETO, leading to an overall relief of line loading in DLR lines, as presented in Fig. 5(b). For example, line 74, a key line connecting the south-west region and Dallas, is heavily loaded in a direction to Dallas (Bus 22 to Bus 94) with an annual average of 84.5% (normalized by line capacity) when operated by DLR baseline. The transition to CHA-UC led to an 6.6%p loading relief. This contributes to an annual conductor health relief of 13.6%, which translates to a 108.8 M\$ reduction in depreciation costs. Here, the commitment decisions took a major role; the annual average commitment on bus 94 increased by 38%, and the commitment on bus 30, adjacent to the east of bus 94, increased by 26.5%, leading to the reduction of power flow to Dallas and stress on line 74.
IV-C Case study II - Impact of Correlated Forecast Errors
@@
IV-C1 Overall Performance Comparison
The costs of the annual simulation are reported on the right-hand side of Table I. Overall, the trend of the three models from case study I was preserved. CHA-UC recorded the lowest annual total cost, 25.41 M\$, contributing to a 34.2%, 24.7% and 28.4% reduction compared to DLR, SLR and QRF, respectively. However, unlike case study I, CHA-UC always incurred the highest day-ahead cost and it recorded the least reserve activation costs. This occurs because only CHA-UC accounts the stochasticity in WP forecasts in the day-ahead decision making.
IV-C2 Impact of Correlated WP and DLR Forecast Error
The effects from Corollary 3 are studied and a physical interpretation is made. Two pairs of wind farms and lines were selected for an analysis on inter-asset correlation. Table II lists the rank correlation of pairs and average line flow using DLR-UC; each are a case of $s_{l}\mathrm{GSF}_{w,l}>0$ and $s_{l}\mathrm{GSF}_{w,l}<0$.
TABLE II: (Case Study II) Pairs of wind farms and lines. GSF was obtained by setting the paired bus as the from-bus.
@@ IV-C2 Impact of Correlated WP and DLR Forecast Error
Figure 6 shows the line-temperature heatmap under baseline DLR. Forecast errors are defined as forecast minus actual, so a larger error indicates greater overestimation. In pair 1, the positive GSF implies that a wind power (WP) shortage reduces line flow magnitude; because the line flow is negative, this means the flow moves away from zero, which stresses the conductor thermal limit. WP forecast errors are often mitigated by DLR forecast errors: DLR underestimation offsets WP shortage by providing additional line capacity in the fourth quadrant, whereas the first quadrant becomes risk-intensive because DLR overestimation amplifies the effect. In contrast, for pair 2, negative WP forecast errors (WP surplus) increase thermal stress, making the second quadrant risk-intensive. Positive correlation concentrates errors in specific quadrants, making pair 1 risk-prone and pair 2 risk-hedging, consistent with Corollary 3.
TABLE III: (Case Study II) Annual average CHA-UC DA line flows with and without scenario correlation
@@
1 -5.79 -5.98 Decreased
2 8.08 7.03 Increased
An experiment regarding correlation proves that the CHA-UC is successfully reflecting the emergent effect. The scenario indices of the pairs were scrambled, changing the marginal minimally, so that the correlation of the scenarios becomes close to 0. This will lead to the neglect of correlation. Separate experiments done with two pairs, in Table III, show that the original (correlation-considering) CHA-UC has kept a more conservative line flow for the risk-prone pair 1 and a less conservative line flow for the risk-hedging pair 2. This asserts that the CHA-UC successfully handles the effect of correlation.
V Conclusion
This paper presents a Conductor Health-Aware Unit Commitment (CHA-UC) framework that explicitly accounts for conductor degradation due to elevated temperature operation (ETO) under DLR. The proposed model integrates a tractable conductor temperature estimation module and a risk-based depreciation cost function, enabling the co-optimization of operating cost and long-term asset health. Case studies demonstrate that CHA-UC reduces total cost by 0.75% and wind power curtailment by 82% compared to SLR operation. It also outperforms a baseline risk-averse approach (QRF-UC), while conventional DLR operation without risk consideration results in higher total costs due to excessive conductor degradation. This work also characterizes how the correlation between wind power and DLR forecast errors can amplify or hedge ETO-induced degradation based on grid conditions. This topology- and flow-dependent effect is explicitly captured by the CHA-UC model. These findings motivate correlation-aware dispatch in operation and highlight the need to account for correlation structures when deploying DLR on wind-connected lines.